The equilibrium constant for the following reaction is 0.18 at a set temperature. Find the equilibrium concentrations if the initial concentration of PCl3 is 0.225 mol/L and the initial concentration of Cl2 is 0.150 mol/L.

PCl3(g) + Cl2(g) ⇌ PCl5(g)




1
Expert's answer
2021-03-05T03:27:44-0500

I:. 0.225............0.150.............0

C:... -x....................-x...................+x

E:..(0.225-x).....(0.150-x)...........x


Kc= [PCl5] / ([PCl3] [Cl2]) = x/[(0.225-x)(0.150-x)] = 0.18

The equation 0.18x2 - 0.9325x + 0.006075= 0 has x=0.0065 as the feasile solution


[PCl5] = 0.0065M

[PCl3] = 0.225 - 0.0065 = 0.2185M

[Cl2] = (0.150 -0.0065) = 0.1435M